OpenStudy (anonymous):

The store only had 5 and ½ bags of cheese left and they needed to spread it evenly on 2 and 2/3 pizzas. How much will each get?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[\huge 5\tfrac{1}{2}\div 2\tfrac{2}{3}\]

OpenStudy (anonymous):

How ? i know its simple but i forgot how

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

turn both in to improper fractions first

OpenStudy (anonymous):

is it cross divide?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

11/2 and 8/3 ?

OpenStudy (anonymous):

i have never heard of "cross divide" \[5\tfrac{1}{2}=\frac{2\times 5+1}{2}=\frac{11}{2}\] right?

OpenStudy (anonymous):

oh yeah you got it

OpenStudy (anonymous):

ok then i divide across?

OpenStudy (anonymous):

\[\frac{11}{2}\div \frac{8}{3}=\frac{11}{2}\times \frac{3}{8}\]

OpenStudy (anonymous):

oh u switch the fraction upside down and multiply so 2 1/33?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok thank you